OpenStudy (anonymous):

sec^2x=4 solve for x a)over the domain of 0_

myininaya (myininaya):

u=secx u^2=4 u=2 u=-2 secx=2 secx=-2

myininaya (myininaya):

cos(pi/3)=1/2 remember so sec(pi/3)=2

myininaya (myininaya):

thats a hint to find your other answers

OpenStudy (anonymous):

thanks